# Probing Green’s Function Zeros by Co-tunneling through Mott Insulators

### v1

@@ Abstract
Quantum tunneling and transport probes are among the most versatile diagnostic tools for gaining insight into the microscopic structure of quantum matter [1, 2, 3, 4]. In the realm of strongly correlated systems [5, 6, 7, 8], experimental breakthroughs along these lines range from single electron control in co-tunneling experiments through quantum dots in the Coulomb blockade regime [9, 10, 11, 12, 13, 14] to probing the fractional statistics of anyonic quasiparticles in topological phases [15, 16, 17, 18]. However, despite this remarkable progress, the complexity of correlated many-body systems still retains a plethora of secrets, thus promising novel physical phenomena yet to be revealed. As an exciting example, Green’s function zeros (GFZ) in Mott insulators [19, 20, 21, 22, 23, 24] have recently been predicted to appear as a shadow of the underlying Bloch band structure [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], i.e., of the poles in the single particle Green’s function (GF) of a quantum material. Owing to Cauchy’s argument principle, such GFZ and poles play a reciprocal role reminiscent of matter and anti-matter in particle physics, and their cancellation at material interfaces has been recently discussed in the context of topological edge modes [33]. Yet, a direct experimental probe of GFZ as a counterpart to well established spectroscopy methods for conventional Bloch bands of GF poles has not been identified.
@@
Figure 1: Cotunneling reveals GFZ Hamiltonian $H^{\text{z}}$: (a) Two spinful auxiliary sites $\{L,R\}$ weakly coupled by tunneling amplitudes $\{\gamma l_{j},\gamma r_{j^{\prime}}\}$ to a Hubbard system in Mott regime. Elastic cotunneling establishes an effective coupling between $L$ and $R$ (red, dashed line). (b) Numerical results (MPS) for the determinate of GF $|\text{det}[G(\omega,k)]|$: SSH-Hubbard model at half-filling with $N=60$ spinful sites, $v=0.75,w=1.0$, $U=10$, analytic prediction of the in-gap zero dispersion of $H^{\text{z}}$ (red, dashed line).
@@
Here, we study co-tunneling through Mott insulators as a novel tool to reveal the shadow band structure of GFZ. As a setting (see Fig. 1 for an illustration), we propose a re-interpretation and generalization of the well-established experimental platform of co-tunneling measurements through quantum dots [9, 35, 36, 37, 38, 14, 39, 40, 41] by replacing a quantum dot in the Coulomb-blockade regime with a many-body system in a Mott phase that may even be engineered as a synthetic quantum material consisting of an array of coupled quantum dots [10, 9, 11]. We pinpoint two fundamental differences between our strongly correlated scenario and a tunneling probe for an uncorrelated band insulator. First, the band structure of GFZ appear in the numerator of the co-tunneling amplitude, contrary to an energy denominator for an uncorrelated insulator. Second, genuine correlation effects modify the zero-bands decisively as compared to the underlying free band structure of the Mott insulator. Using the striking example of topological distinction between zero-bands and the corresponding pole-bands, we demonstrate how these subtle fingerprints of many-body correlations can be directly observed through the co-tunneling amplitude.
@@
– To describe how to probe a solid-state system spectroscopically, we use the formalism of Green’s functions. At the one-particle level, this allows to quantify the response of the system upon adding or removing an electron, taking into account correlations effects. The single-particle electronic Green’s function in the Källén-Lehmann representation at zero temperature is explicitly dependent on the spectrum of excitations of the system [42, 43]:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)$ $=\sum_{\nu}\Bigg{[}\frac{\langle GS|c_{j,\sigma}|E_{\nu}\rangle\langle E_{\nu}|c_{j^{\prime},\sigma^{\prime}}^{\dagger}|GS\rangle}{\omega+E_{GS}-E_{\nu}}$
$+\frac{\langle GS|c_{j^{\prime},\sigma^{\prime}}^{\dagger}|E_{\nu}\rangle\langle E_{\nu}|c_{j,\sigma}|GS\rangle}{\omega+E_{\nu}-E_{GS}}\Bigg{]}\text{.}$ (1)
@@
where $c^{(\dagger)}_{j,\sigma}$ denote the annihilation and creation operators for site $j$ and spin $\sigma$. In the non-interacting limit, the Hamiltonian and $G(\omega)$ are readily diagonalizable in a single-particle basis. The determination of the spectral and geometrical properties of the system then amounts to the study of the poles of $G(\omega)$, each of which represents an infinitely long-lived single-particle eigenmode. However, in presence of correlations, a large amount of matrix elements in the numerator become nonzero, reflecting the multitude of scattering events promoted by the many-body interaction term. As a result, the spectral weight distribution is broadened in the frequency-momentum plane. There is, however, a remarkable consequence of this scenario: in the spectral distribution, a set of isolated Green’s function zeros (GFZ) can emerge. Their presence has been linked to peculiar thermodynamic [44] and topological properties [25, 26, 33, 34, 27, 28, 29, 30, 31, 32], and the study of their dispersion is hence an intriguing, if nontrivial endeavor.
@@
The GFZ are defined as zero eigenvalues of $G$, or equivalently as the points at which its determinant vanishes [25]. In the noninteracting limit, there are no isolated zeros. In the opposite case, where the system is deep in the Mott phase, two incoherent Hubbard bands form at high frequency, which contain the totality of the poles of the Green’s function, and a large number of zeros. However, inside the hard Mott gap between the two Hubbard bands no poles are present. By contrast, GFZ can and do manifest at low energy, irrespective of the system size. The focus of this work lies on how to detect these isolated in-gap GFZ which are entirely responsible for the low energy properties of the GF of Mott insulators [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
As a first step in our discussion, we derive a remarkably simple and intuitive expression for $G$ at low energies ($\omega\ll U$). Our non-interacting Hamiltonian $H_{0}=\sum h_{i\sigma;j\sigma^{\prime}}c^{\dagger}_{i,\sigma}c_{j,\sigma^{\prime}}$ consists of a linear combination of on-site terms, which we denote by $M$, and non-local hopping terms ($W$). The interaction term is of the Hubbard type $UH_{\text{int}}=2U\sum_{j}(n_{j,\downarrow}-\frac{1}{2})(n_{j,\uparrow}-\frac{1}{2})$, where $j$ labels the lattice site and $\sigma=\uparrow,\downarrow$ is the spin. The two competing energy scales are the bandwidth of $H_{0}$, $2D$ and the local Coulomb interaction strength $U$. In the half-filled Mott insulator, $U\gg D$. For convenience, we introduce the rescaled non-interacting Hamiltonian term $H^{\prime}_{0}=\frac{1}{D}H_{0}$, as well as the small parameter $\phi=\frac{D}{U}$. The Hamiltonian can therefore be rewritten as
@@
In this form, $H_{\text{H}}$ naturally suggests a perturbative analysis in $\phi$: indeed, by applying degenerate perturbation theory [45] up to first order in $\phi$ we determine the perturbed eigenenergies eigenenergies $E_{\nu}^{P}$ and eigenstates $|E_{\nu}^{P}\rangle$, which we can then conveniently insert in Eq. (1) to obtain the perturbed Mott Green’s function. It is finally a matter of collecting all the lowest terms in $\phi$, to come to a key expression for the Green’s function which is valid at low energies $\omega\ll U$  [^1] Due to expansions of $\frac{1}{\omega-\omega_{0}}$ terms, our approximation breaks down at large energies $\omega\approx U$.:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)\underset{\omega,D\ll U}{\approx}\frac{1}{U^{2}}\left[-\omega\mathbb{I}+H^{\text{z}}\right]_{j,\sigma;j^{\prime},\sigma^{\prime}}+\mathcal{O}(\phi^{3}).$ (3)
This expression is intuitively consistent with the strongly suppressed in-gap spectral distribution of a Mott insulator. More than that, it immediately and clearly highlights the presence, inside of the Mott gap, of points where the determinant of the Green’s function is exactly zero. These are the roots of the expression $H^{z}-\omega\mathbb{I}$. In other words, and inline with [33], we find that the low-energy isolated zeros of a Mott insulator are described by an effective Hamiltonian $H^{\text{z}}$. The detailed derivation of Eq. (3) can be found in the supplemental material [47].
@@
For strong Coulomb repulsion $U$ and at half-filling, the Hubbard model at low energies is well described by an effective spin-$1/2$ model [7, 8, 5, 6]. We can therefore rewrite the lowest-order terms in $G$ in the spin basis, thereby obtaining an expression linking the approximated Green’s function to spin-spin correlators, as derived in detail in [47]. For a $SU(2)$-symmetric groundstate and bare Hamiltonian $(H_{0})_{j,\sigma;j^{\prime},\sigma^{\prime}}=(M+W)_{jj^{\prime}}\delta_{\sigma,\sigma^{\prime}}$, we find a remarkably elegant expression for the effective Hamiltonian of the GFZ:
@@
where $j,j^{\prime}$ are lattice sites. Here, expectation values are evaluated with respect to the groundstate eigenvector $|GS^{0}\rangle$ of the corresponding effective spin-$1/2$ model  [^2] That is the correct groundstate towards the unperturbed limit and follows from degenerated perturbation theory., and are hence accessible for a large variety of Hubbard models. A general discussion of $SU(2)$-breaking bare Hamiltonians is presented in [47], while we will restrict ourselves in the following to $SU(2)$-symmetric systems. The effective zeros Hamiltonian can then be decomposed in two terms, in a striking analogy with the on-site and hopping terms of the bare Hamiltonian $H_{0}$: (i) the on-site terms of the non-interacting Hamiltonian, described by $M$, are carried over in the dispersion of the zeros without renormalization, while (ii) the hopping terms, described by $W$, are re-scaled by the spin correlation function $\langle\overrightarrow{S}_{j}\cdot\overrightarrow{S}_{j^{\prime}}\rangle$  [^3] The correlation value is evaluated in the limit towards, not at, the atomic limit, making it generally non-trivial.. This follows naturally from the combination of our perturbative approach around the atomic limit and the nature of competing energy scales in the system: any quadratic on-site term commutes with $H$ in the atomic limit, and the perturbed energy is just shifted by $M$ without renormalization. On the contrary, hopping and interaction strength are in direct competition: treating the first order perturbatively results in a lowest-order correction to the atomic limit Hamiltonian which famously depends on the spin-spin correlation function, as in the derivation of the superexchange [50, 51, 52, 8].
@@
– As clearly evidenced by Eqs. (3) and (4), the knowledge of the zeros dispersion is sufficient to determine the spectral properties of the Mott system at low frequency. Obtaining a clear hint of the zeros’ behavior from experimental observations is instead less straightforward since any spectral measurement (e.g. via photoemission) would just give confirmation of the presence of a gap, without being able to resolve the elusive zeros within. We propose a novel approach to the problem which is, at the same time, surprisingly simple to implement: instead of directly performing spectroscopic measurements on the Mott insulator, we employ it as the potential barrier in a co-tunneling experiment, and study the effective coupling between the leads. We illustrate the setup via a minimal example, which is a convenient adaption of already existing quantum dot experimental setups [36, 11, 9, 10]. This is sketched in Fig.1 (a) and described by
$H_{\text{tot}}=H_{\text{H}}+\gamma C+H_{\text{D}}$ (5)
The system consists of two quantum dots, henceforth L and R, and a central tunneling medium. The Hamiltonian of the quantum dots, $H_{\text{D}}$, is that of a Hubbard atom with Coulomb repulsion $2U$, while the central element is a spin-isotropic Hubbard model in the Mott phase, described by the previously discussed Eq. (2). The chemical potential of the quantum dots is increased by $U$ in relation to the central Hubbard model, and independently adjustable by $\mu_{\text{L}},\mu_{\text{R}}$. The system-dot coupling $C$ is given by arbitrary hopping amplitudes $\{l_{i}\}$ and $\{r_{i}\}$ between the $i$-th site of the central Mott insulator and the L and R quantum dots. For simplicity, we will from now on omit the spin index $\sigma$, and restrict ourselves to the case where $l_{i}=\delta_{i,j}$ and $r_{i}=\delta_{i,j^{\prime}}$ for two sites $j\neq j^{\prime}$. The full derivation for the general case with arbitrary hopping array is provided in [47]. In a co-tunneling experiment, only the leads are directly probed. This amounts to tracing away the degrees of freedom associated to the central system, and only considering the overall propagation amplitude of excitation leaving (entering) the two leads via the quantum dots. These processes are, in our proposed setting, exactly described by the one-particle Green’s function of the central Mott insulator. If we are interested in the low-frequency excitations of the system, the picture can be simplified further, making use of the previously derived approximation for the Green’s function. In short, by applying two Schrieffer-Wolff transformations [53, 54, 55] in the limit of weak system-dot coupling $\gamma\ll D,U$, and assuming a large interaction $U$, we derive an effective spin-isotropic low energy Hamiltonian $\bar{h}$. It describes the dynamics of a single particle co-tunneling between the left and right quantum dots through the half-filled Mott-insulator:
@@
where $\bar{\mu}$ is an “average” chemical potential $\frac{1}{2}\left(\mu_{\text{L}}+\mu_{\text{R}}\right)$. The presence of the central Mott insulator provides a renormalization to the on-site energies, given by $u_{L/R}\sim\mathcal{O}(\frac{\gamma^{2}}{U})$, which can be conveniently reabsorbed by tuning the quantum dot chemical potentials. The most relevant parameter in Eq. (6) is the amplitude
$t_{\text{co}}(\bar{\mu})=\gamma^{2}G_{j,j^{\prime}}(\bar{\mu}),$ (7)
which describes the co-tunneling mediated coupling between the left and right quantum dot via the central Mott insulator. We immediately to notice how, through Eq. (3), the co-tunneling amplitude $t_{\text{co}}$ is directly related to the GFZ-Hamiltonian $H^{\text{z}}$ of Eq.(4) via $t_{\text{co}}=\frac{\gamma^{2}}{U^{2}}\left(H^{\text{z}}\right)_{j,j^{\prime}}$. This clarifies the correspondence between the Mott Green’s function and the tunneling amplitude: instead of directly probing the features of the vanishing spectral function inside the Mott gap, a nearly impossible task, we are now recasting the same information into what is effectively a potential barrier for the evanescent wave of the co-tunneling experiment. Its features, and in particular the spectroscopically undetectable zeros, have a macroscopic effect on the measured occupations of the leads. As a further benefit, we can now take advantage of the well established theoretical tools and experimental techniques employed in the study of quantum tunneling, such as charge stability diagrams [56, 57, 58, 11, 59]and time-resolved charge detection of single tunneling events [60, 61, 62, 36].
@@
– We now provide numerical evidence of our key findings by considering the paradigmatic case of the one-dimensional Su-Schrieffer-Heeger (SSH)-Hubbard model[33, 63], consisting of a bipartite chain with alternating hoppings $v$ and $w$, and arbitrary on-site potentials. We numerically solve the Hamiltonian of the co-tunneling setup by using a combination of Density Matrix Renormalization Group (DMRG) [64, 65], time-dependent variational principle (TDVP) [66, 67, 68] and exact diagonalization (ED) methods, by first computing the real time GF and then transforming it into the frequency domain, thus gaining access to the GFZ.
We first prove that our analytical formula for the zeros Hamiltonian is an adequate approximation of the in-gap spectral features of a Mott insulator. Fig.1 (b) shows a comparison between the determinant of the Green’s function of the SSH-Hubbard model, obtained via DMRG, and the approximated formula for the zeros effective Hamiltonian of Eq. (4). We note that the isolated zeros (dark purple regions of the color plot), which are formed in the Mott gap, are in perfect agreement with the eigenvalues of $H^{\text{z}}$ already at the intermediate value of $\phi\approx 1/10$.
Figure 2: Co-tunneling amplitude $t_{\text{co}}$ as function of bandwidth $D$(fixed $U=2000$, $N=16$, $\gamma=20$,$l_{i}=\delta_{i,2}$,$r_{i}=\delta_{i,3}$, $v=0.9w$, PBC). SSH-Hubbard model (solid,red) shows linear $D$ dependence within the the Mott regime $\frac{D}{U}<1$, due to energy nominator in GF Eq.(3). Non-interacting insulator (solid,grey) with energy denominator (parameters as above but $U=0$). Analytical results, Eq.(7) (dashed).

### v2

@@ Abstract
Quantum tunneling and transport probes are among the most versatile diagnostic tools for gaining insight into the microscopic structure of quantum matter [1, 2, 3, 4]. In the realm of strongly correlated systems [5, 6, 7, 8], experimental breakthroughs along these lines range from single electron control in co-tunneling experiments through quantum dots in the Coulomb blockade regime [9, 10, 11, 12, 13, 14] to probing the fractional statistics of anyonic quasiparticles in topological phases [15, 16, 17, 18]. However, despite this remarkable progress, the complexity of correlated many-body systems still retains a plethora of secrets, thus promising novel physical phenomena yet to be revealed. As an exciting example, Green’s function zeros (GFZ) in Mott insulators [19, 20, 21, 22, 23, 24] have recently been predicted to appear as a shadow of the underlying Bloch band structure [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], i.e., of the poles in the single particle Green’s function (GF) of a quantum material. Owing to Cauchy’s argument principle, such GFZ and poles play a reciprocal role reminiscent of matter and anti-matter in particle physics, and their cancellation at material interfaces has been recently discussed in the context of topological edge modes [33]. Despite recent progress regarding the possibility of detecting GFZ either by exploiting zero-pole annihilation mechanisms [33, 34, 35, 36] or through transport signatures [27], it is fair to say that a direct experimental probe of GFZ has largely remained elusive.
@@
Figure 1: Co-tunneling reveals GFZ-Hamiltonian $H^{\text{z}}$: (a) Two spinful auxiliary sites (grey, $\{L,R\}$), weakly coupled by tunneling amplitudes $\{\gamma l_{j},\gamma r_{j^{\prime}}\}$ to a Hubbard system in Mott regime (violet). Elastic co-tunneling establishes an effective coupling between $L$ and $R$ (red, dashed line). (b) Numerical results (MPS) on the GF-determinant $|\text{det}[G(\omega,k)]|$: SSH-Hubbard model at half-filling with $N=60$ spinful sites, $v=0.75,w=1.0$, $U=10$ and analytic prediction of the in-gap zero dispersion of $H^{\text{z}}$ (red, dashed line).
@@
Here, we study co-tunneling through Mott insulators as a sharp diagnostic tool to fully reveal the shadow band structure of GFZ. As a setting (see Fig. 1 for an illustration), we propose a re-interpretation and generalization of the well-established experimental platform of co-tunneling measurements through quantum dots [9, 37, 38, 39, 40, 14, 41, 42, 43] by replacing a quantum dot in the Coulomb-blockade regime with a many-body system in a Mott phase, which may even be engineered as a synthetic quantum material consisting of an array of coupled quantum dots [10, 9, 11, 44]. We pinpoint two fundamental differences between our strongly correlated scenario and a tunneling probe for an uncorrelated band insulator. First, the band structure of GFZ appears in the numerator of the co-tunneling amplitude, contrary to an energy denominator for an uncorrelated insulator. Second, genuine correlation effects decisively modify the zero-bands compared to the underlying free band structure of the Mott insulator. Using the striking example of topological distinction between zero-bands and the corresponding pole-bands, we demonstrate how these subtle fingerprints of many-body correlations can be directly observed through the co-tunneling amplitude in mesoscopic quantum transport experiments.
@@
– To describe how to probe a solid-state system spectroscopically, we use the formalism of Green’s functions. At the one-particle level, this allows to quantify the response of the system upon adding or removing an electron, taking into account correlation effects. The single-particle electronic Green’s function in the Källén-Lehmann representation at zero temperature is explicitly dependent on the spectrum of excitations of the system [45, 46]:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)$ $=\Bigl{<}GS\Big{|}\Bigl{[}c_{j,\sigma}(\omega+E_{GS}-H_{\text{H}})^{-1}c_{j^{\prime},\sigma^{\prime}}^{\dagger}$
$~~+c_{j^{\prime},\sigma^{\prime}}^{\dagger}(\omega-E_{GS}+H_{\text{H}})^{-1}c_{j,\sigma}\Bigr{]}\Big{|}GS\Bigr{>}\text{,}$ (1)
@@
where $c^{(\dagger)}_{j,\sigma}$ denotes the annihilation and creation operators for site $j$ and spin $\sigma$. In the non-interacting limit, the Hamiltonian and $G(\omega)$ are readily diagonalizable in a single-particle basis. The determination of the spectral and geometrical properties of the system then amounts to the study of the poles of $G(\omega)$, each of which represents an infinitely long-lived single-particle eigenmode. However, in presence of correlations, a large amount of amplitudes in the numerator become nonzero, reflecting the multitude of scattering events promoted by the many-body interaction term. As a result, the spectral weight distribution is broadened in the frequency-momentum plane. There is, however, a remarkable consequence of this scenario: in the spectral distribution, a set of isolated Green’s function zeros (GFZ) can emerge. Their presence has been linked to peculiar thermodynamic [47] and topological properties [25, 26, 33, 34, 27, 28, 29, 30, 31, 32], and the study of their dispersion is therefore an intriguing, if nontrivial endeavor.
@@
The GFZ are defined as zero eigenvalues of $G$, or equivalently as the points at which its determinant vanishes [25]. In the noninteracting limit, there are no isolated zeros. In the opposite case, where the system is deep in the Mott phase, two incoherent Hubbard bands form at high frequency, which contain the totality of the poles of the Green’s function, and a large number of zeros. However, inside the hard Mott gap between the two Hubbard bands, no poles are present. By contrast, GFZ can, and do, manifest at low energy, irrespective of the system size. The focus of this work lies on how to detect these isolated in-gap GFZ, which are entirely responsible for the low energy properties of the GF of Mott insulators [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
As a first step in our discussion, we derive a remarkably simple and intuitive expression for $G$ at low energies ($\omega\ll U$) deep in the Mott regime at half-filling, and with a paramagnetic groundstate. Our non-interacting Hamiltonian $H_{0}=\sum h_{j;j^{\prime}}c^{\dagger}_{j,\sigma}c_{j^{\prime},\sigma}$ consists of a linear combination of on-site terms, which we denote by $M$, and non-local hopping terms ($W$). The interaction term is of the Hubbard type $UH_{\text{int}}=2U\sum_{j}(n_{j,\downarrow}-\frac{1}{2})(n_{j,\uparrow}-\frac{1}{2})$, where $j$ labels the lattice site and $\sigma=\uparrow,\downarrow$ is the spin. The two competing energy scales are the bandwidth of $H_{0}$, $2D$, and the local Coulomb interaction strength $U$, where $U\gg D$ for a Mott insulator. For convenience, we introduce the rescaled non-interacting Hamiltonian term $H^{\prime}_{0}=\frac{1}{D}H_{0}$, as well as the small parameter $\phi=\frac{D}{U}$. The Hamiltonian can therefore be rewritten as
@@
In this form, $H_{\text{H}}$ naturally suggests a perturbative analysis in $\phi$: indeed, by applying degenerate perturbation theory [48] up to first order in $\phi$ we determine its perturbed eigenenergies $E_{\nu}^{P}$ and eigenstates $|E_{\nu}^{P}\rangle$, which we can then conveniently insert in Eq. (1) to obtain the perturbed Mott Green’s function. It is finally a matter of collecting all the lowest terms in $\phi$ and expanding the denominator at low energies, to arrive at a key expression for the Green’s function, which is valid for $\phi\ll 1$ and $\omega\ll U$  [^1] Due to expansions of $\frac{1}{\omega-\omega_{0}}$ terms, our approximation breaks down at large energies $\omega\approx U$.:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)\,\,{\approx}\,\,\frac{1}{U^{2}}\left[-\omega\mathbb{I}+H^{\text{z}}\right]_{j,\sigma;j^{\prime},\sigma^{\prime}}+\mathcal{O}(\phi^{3}).$ (3)
This expression is intuitively consistent with the strongly suppressed in-gap spectral distribution of a Mott insulator. More than that, it immediately and clearly highlights the presence, inside of the Mott gap, of points where the determinant of the Green’s function is exactly zero. These are the roots of the expression $H^{z}-\omega\mathbb{I}$. In other words, and inline with [33], we find that the low-energy isolated zeros of a Mott insulator are described by an effective Hamiltonian $H^{\text{z}}$. The detailed derivation of Eq. (3) can be found in the supplemental material [50], including a general discussion of $SU(2)$-breaking bare Hamiltonians and groundstates.
@@
For strong Coulomb repulsion $U$ and at half-filling, the Hubbard model at low energies is well described by an effective spin-$1/2$ model [7, 8, 5, 6]. We can therefore rewrite the lowest-order terms in $G$ in the spin basis, thereby obtaining an expression linking the approximated Green’s function to spin-spin correlators, as derived in detail in [50]. For the bare Hamiltonian $h_{j;j^{\prime}}=(M+W)_{jj^{\prime}}$, we find a remarkably elegant expression for the effective Hamiltonian of the GFZ:
@@
where $j,j^{\prime}$ are lattice sites. Here, expectation values are evaluated with respect to the groundstate eigenvector $|GS^{0}\rangle$ of the corresponding effective spin-$1/2$ model  [^2] That is the correct groundstate towards the unperturbed limit and follows from degenerated perturbation theory., and are hence accessible for a large variety of Hubbard models. The effective zeros Hamiltonian can then be decomposed into two terms, in a striking analogy with the on-site and hopping terms of the bare Hamiltonian $H_{0}$: (i) the on-site terms of the non-interacting Hamiltonian, described by $M$, are carried over in the dispersion of the zeros without renormalization, while (ii) the hopping terms, described by $W$, are re-scaled by the spin correlation function $\langle\overrightarrow{S}_{j}\cdot\overrightarrow{S}_{j^{\prime}}\rangle$  [^3] The correlation value is evaluated in the limit towards, not at, the atomic limit, making it generally non-trivial.. This follows naturally from the combination of our perturbative approach around the atomic limit, and the nature of competing energy scales in the system: any quadratic on-site term commutes with $H$ in the atomic limit, and the perturbed energy is just shifted by $M$ without renormalization. On the contrary, hopping and interaction strength are in direct competition: treating the first order perturbatively results in a lowest-order correction to the atomic limit Hamiltonian, which famously depends on the spin-spin correlation function, as in the derivation of the superexchange [57, 58, 59, 8]. These antiferromagnetic correlations characterize paramagnetic Mott insulators beyond mean field and are responsible for the sign flip in the nearest-neighbour hopping in Eq. (4).
@@
– As clearly evidenced by Eqs. (3) and (4), the knowledge of the zeros dispersion is sufficient to determine the spectral properties of the Mott system at low frequency. Obtaining a clear hint of the zeros behavior from experimental observations is instead less straightforward since any spectral measurement (e.g. via photoemission) would just confirm the presence of a gap, without being able to resolve the elusive zeros within. We propose a novel approach to the problem, which is, at the same time, surprisingly simple to implement: instead of directly performing spectroscopic measurements on the Mott insulator, we employ it as the potential barrier in a co-tunneling experiment, and study the effective coupling between the leads. We illustrate the setup via a minimal example, which is a convenient adaption of already existing quantum dot experimental setups [38, 11, 9, 10]. This is sketched in Fig.1 (a) and described by $H_{\text{tot}}=H_{\text{H}}+\gamma C+H_{\text{D}}$. The system consists of two quantum dots, henceforth L and R, and a central tunneling medium. The Hamiltonian of the quantum dots, $H_{\text{D}}$, is that of a Hubbard atom with Coulomb repulsion $2U$. In order to reveal fingerprints of the GFZ, the central element is a spin-isotropic Hubbard model in the Mott phase, described by the previously discussed Eq. (2). The chemical potential of the quantum dots is increased by $U$ in relation to the central Hubbard model, and independently adjustable by $\mu_{\text{L}},\mu_{\text{R}}$. The system-dot coupling $C$ is given by arbitrary hopping amplitudes $\{l_{i}\}$ and $\{r_{i}\}$ between the $i$-th site of the central Mott insulator and the L and R quantum dots. For simplicity, we will, from now on, omit the spin index $\sigma$, and restrict ourselves to the case where $l_{i}=\delta_{i,j}$ and $r_{i}=\delta_{i,j^{\prime}}$ for two different sites $j\neq j^{\prime}$. The full derivation for the general case with arbitrary hopping array is provided in [50].
In a co-tunneling experiment, only the leads are directly probed. This amounts to tracing out the degrees of freedom associated to the central system, and only considering the overall propagation amplitude of excitation leaving (entering) the two leads via the quantum dots. These processes are, in our proposed setting, exactly described by the one-particle Green’s function of the central tunneling medium. In short, by applying a Schrieffer-Wolff transformation [60, 61, 62] in the limit of weak system-dot coupling $\gamma$ with small dot potentials $\mu_{\text{L}/\text{R}}$  [^4] More precise: the system-dot coupling $\gamma$ and the dot potentials $\mu_{\text{L}/\text{R}}$ need to be smaller than the first excitation gap above the groundstate of the central barrier. , and assuming a large lead-dot interaction $\sim U$, we derive an effective spin-isotropic Hamiltonian $\bar{h}$ in the low-energy subspace. It describes the dynamics of a single particle co-tunneling between the left and right quantum dots through the central medium:
@@
where $\bar{\mu}$ is an ”average” chemical potential $\frac{1}{2}\left(\mu_{\text{L}}+\mu_{\text{R}}\right)$. The presence of the central tunneling medium gives rise to a renormalization to the on-site energies, given by $u_{L/R}\sim\mathcal{O}(\gamma^{2})$, which can be conveniently reabsorbed by tuning the quantum dot chemical potentials. The most relevant quantity in Eq. (5) is the co-tunneling amplitude
$t_{\text{co}}(\bar{\mu},l_{j},r_{j^{\prime}})=\gamma^{2}G_{j,j^{\prime}}(\bar{\mu}),$ (6)
which describes the coupling between the left and right quantum dot mediated via the central barrier (see Fig. 1 for an illustration). Since we are interested in a Mott insulator at half-filling as the tunneling barrier, we can make use of the previously derived approximation for the Green’s function in Eq. (3). We immediately notice how, through Eq. (3), the co-tunneling amplitude $t_{\text{co}}$ is directly related to the GFZ-Hamiltonian $H^{\text{z}}$ of Eq. (4). Explicitly, for $l_{j}$ and $r_{j^{\prime}}$ (cf. Fig. 1) chosen such that $j\neq j^{\prime}$ (see [50] for general tunnel-couplings $l,r$):
$t_{\text{co}}=\frac{\gamma^{2}}{U^{2}}\left(H^{\text{z}}\right)_{j,j^{\prime}}$ (7)
As a central result, Eq. 7 clarifies the correspondence between the Mott Green’s function and the co-tunneling amplitude: instead of directly probing the features of the vanishing spectral function inside the Mott gap, a nearly impossible task, we are now recasting the same information into what is effectively a potential barrier for the evanescent wave of the co-tunneling experiment. Its features, and in particular the spectroscopically undetectable zeros, have a macroscopic effect on the measured occupations of the leads. As a further benefit, we can now take advantage of the well-established theoretical tools and experimental techniques employed in the study of quantum tunneling, such as charge stability diagrams [64, 65, 66, 11, 67] and time-resolved charge detection of single tunneling events [68, 69, 70, 71, 72, 73, 38].
@@
– We now provide numerical evidence of our key findings by considering the paradigmatic case of the one-dimensional Su-Schrieffer-Heeger (SSH)-Hubbard model[33, 74], consisting of a bipartite chain with alternating hoppings $v$ and $w$. We numerically solve the Hamiltonian of the co-tunneling setup by using Density Matrix Renormalization Group (DMRG) [75, 76] in combination with the time-dependent variational principle (TDVP) [77, 78, 79] and exact diagonalization (ED) methods. We first compute the real-time GF and then transform it into the frequency domain, thereby gaining access to the GFZ.
We first prove that our analytical formula for the zeros Hamiltonian is an adequate approximation of the in-gap spectral features of a Mott insulator. Fig.1 (b) shows a comparison between the determinant of the Green’s function of the SSH-Hubbard model, obtained via DMRG, and the approximated formula for the GFZ Hamiltonian of Eq. (4). We note that the isolated zeros (dark purple regions of the color plot), which appear in the Mott gap, are in perfect agreement with the eigenvalues of $H^{\text{z}}$, already at the intermediate value of $\phi\approx 1/10$. In the supplemental material [50], we provide a more detailed quantitative analysis.
Figure 2: Co-tunneling amplitude $t_{\text{co}}$ as function of bandwidth $D$ (fixed $U=2000$, $N=16$, $\gamma=20$$l_{i}=\delta_{i,2}$$r_{i}=\delta_{i,3}$, $v=0.9w$, PBC). SSH-Hubbard model (solid,red) shows linear $D$ dependence within the the Mott regime $\frac{D}{U}<1$, due to energy nominator in GF Eq. (3). Non-interacting insulator (solid,grey) with energy denominator (parameters as above but $U=0$). Analytical results, Eq.(6) (dashed).